OpenStudy (anonymous):

What is the specific heat of an unknown substance if 2000 J of energy are required to raise the temperature of 4 grams of the substance 5 degrees Celsius?

OpenStudy (anonymous):

what is specific heat?

OpenStudy (anonymous):

4.2J=1cal

OpenStudy (anonymous):

1g, 1degree

OpenStudy (anonymous):

something

OpenStudy (michele_laino):

here is the formula of the specific heat: \[\Large c = \frac{Q}{{m\Delta T}} = \frac{{2000}}{{4 \cdot {{10}^{ - 3}} \cdot 5}} = ...\frac{{joules}}{{Kg \cdot kelvin}}\]

OpenStudy (michele_laino):

please note that, the quantity \(\Delta T\) can be measure with celsius degrees or with Kelvin degrees, since it is a \(difference\) bewtween two temperatures

OpenStudy (michele_laino):

measured*